# InComeS: Integrating Compression and Selection Mechanisms into LLMs for Efficient Model Editing

### v2

@@ C.1 Datasets
The dataset MQuAKE [49] (Multi-hop Question Answering for Knowledge Editing) is constructed based on Wikidata and contains question answering instances that require 2-hop, 3-hop, and 4-hop reasoning. In the experiment, we use the latest version of the dataset 999”MQuAKE-CF-3k-v2.json” in https://github.com/princeton-nlp/MQuAKE, which fixes the knowledge conflict problem for the old version multi-edit subset, and report the accuracy for each query.
@@
DUNE [1] is a benchmark designed for edits in natural language form. It evaluates the model’s capability of conducting natural language edits through four aspects: scientific reasoning, arithmetic reasoning, new information, and debiasing. As illustrated in Table 2 of [1], the arithmetic reasoning edits do not follow natural language form as other subsets do and cannot represent a complete piece of instruction, therefore, we do not include it in our experiment.
WikiDatacounterfact
The $\text{WikiData}_{counterfact}$ [2, 46] collects triplets from top-viewed pages from Wikipedia and contains portability (ripple-effect [2]) instances to test whether the output to the input relevant to the edits is changed as well. Specifically, the portability evaluates the post-edited model from three aspects, including logical generalization, subject aliasing, and reasoning.
@@
The extended version of ZsRE [46, 43] is constructed based on the original ZsRE [20], which is a dataset that focuses on the QA task. The extended version introduces a portability test [43], including inverse relation, one-hop reasoning, and subject aliasing.
@@
COUNTERFACT [25] is a dataset that concentrates on counterfactual information, which typically receives a lower prediction score than accurate facts. It constructs out-of-scope data by substituting the subject entity with a comparable description that has the same predicate.
C.2 Evaluation metrics
This section explains the evaluation metrics used in the extended ZsRE[43, 46] and $\text{Wiki}_{counterfact}$ [2, 46]. Generally, they adopt four metrics: reliability, generality, portability, and locality. Given an initial base model $f_{\theta}$, a post-edit model $f_{\theta^{\prime}}$, and a set of edit instances $(x_{t},y_{t})\in\{(x_{t},y_{t})\}$, the reliability is computed as the average accuracy of the edit cases:
@@ C.2 Evaluation metrics
For the extended ZsRE [43, 46] and $\text{Wiki}_{counterfact}$ [2, 46], we follow the setting in the original paper and combine reliability and generality to the Edit Success rate.
C.3 Baseline implementation details
Unless otherwise specified, the baselines are implemented by using the EasyEdit framework [39].
@@ C.3 Baseline implementation details
We follow the procedure implemented in previous work [24, 25, 43, 46] to fine-tune a specific layer from the model. We select layer 13 for Llama-3.2-1B and layer 27 for Qwen2.5-7B. For both models, we adopt the learning rate of $5e^{-4}$ and the number of optimization steps 25.
@@
For both models, we use LoRA [16] to update the query and key projection matrix of the models, with rank set to 8, $\alpha$ set to 32, the dropout rate 0.1, and the learning rate $5e^{-3}$. The number of updating steps is set to 70 for Llama-3.2-1B and 60 for Qwen2.5-7B.
@@
ROME [24] treats the FFN part of the LLMs as a key-value association and updates a pre-located layer by directly inserting an optimized key-value pair. We update the layer 5 for both Llama-3.2-1B and Qwen2.5-7B, and adopt 25 optimization steps for Llama-3.2-1B and 20 optimization steps for Qwen2.5-7B, with both learning rate $5e^{-5}$.
@@
R-ROME [10] is another version of ROME [24] with modified code implementation. We use the same hyperparameters as ROME.
@@
KN [3] hypothesize that factual knowledge is stored in FFN memories and expressed by knowledge neurons. For both models, we use the threshold of 0.2 for knowledge attribution scores and 0.4 for the threshold of the prompts sharing percentage.
@@
GRACE [12] adopts a discrete codebook to memorize the edits as key-value pairs. We set the location of the codebook layer 13 and 18 for Llama-3.2-1B and Qwen2.5-7B, respectively. Surprisingly, the $\epsilon$ value used in the original paper (1-3) seems insufficient for the complex editing experiments in this paper. Therefore, we increase it to 50. The number of optimization steps for the value vector is set to 100.
@@
IKE [48] maintains an explicit memory for edits and retrieves them via K-nearest neighbors. The retrieved edits are then used to construct demonstrations, which are then prefixed to the input to edit the behavior. In the experiments, we set $K=16$.
@@
MEND [28] trains an additional meta-network to predict a new rank-one update to the input gradient. In this paper, we train each model using ZsRE [20] and COUNTERFACT [25], and adopt the ZsRE-trained model to MQuAKE and the extended ZsRE and COUNTERFACT-trained model for $\text{Wiki}_{counterfact}$.
@@
SERAC [29] employs an explicit edit-instance memory, an additional trained scope classifier, and a trained counterfactual model. The scope classifier is responsible for determining whether an input is relevant to the edits in the memory. The input is fed to the counterfactual model once the input is deemed as relevant to memorized edits and the original model otherwise. We use the distilbert-base-cased [38] model as the scope classifier, and train it using the training set of ZsRE [20] and COUNTERFACT [25] from EasyEdit [39]. Following [1], we use instruction-tuned models (Llama-3.2-1B-Instruct [^10] https://huggingface.co/meta-llama/Llama-3.2-1B-Instruct and Qwen2.5-0.5B [^11] https://huggingface.co/Qwen/Qwen2.5-0.5B-Instruct) for the counterfactual model.
@@
MEMIT [25] is the extension of ROME [24] that supports a batch of edits at a time. Unlike ROME, which only updates a single pre-located layer, MEMIT spreads the update to a set of identified layers. We apply changes to layers 4-8 for both models. Following the settings in the original paper, we set $\lambda$, the hyperparameter that balances the weighting of new and old associations, to $1.5\times 10^{4}$.
@@
EMMET [11] is an unification of ROME [24] and MEMIT [25]. Similar to ROME, we edit the layer 5 for both models, and set $\lambda=1e^{5}$.
@@
We adopt bge-base-en-v1.5 [41] as our retriever. For batch editing, we treat the corresponding batch number of edits as our corpus, and retrieve the 10 most relevant edits for each testing query.
@@
Appendix D Further Analysis
D.1 RAG vs. InComeS
@@ Appendix D Further Analysis
Method Model Single Editing Batch Editing 2-edits 3-edits 4-edits 2-edits 3-edits 4-edits Base Llama-3.2-1B 41.79 43.51 31.58 41.79 43.51 31.58 RAG 59.23 59.00 51.63 40.65 46.78 43.58 ICL 59.23 59.00 51.63 50.06 49.86 42.37 InComeS 71.19 72.17 72.62 53.93 52.79 52.73 Base Qwen2.5-7B 44.08 44.14 30.62 44.08 44.14 30.62 RAG 69.76 76.91 74.54 42.91 46.37 46.15 ICL 69.76 76.91 74.54 53.53 50.54 44.77 InComeS 66.46 71.24 76.54 55.13 53.48 47.91
Table 8: Results for RAG on MQuAKE [49].
D.2 Side effect
We present a side effect analysis of our method in this section (Table 9). We test the editing side effect under three different numbers of edits (0, 0.1k, 1k) on MMLU [13] benchmark, which consists of 57 tasks across 4 domains, namely Social science, Humanities, STEM, and others. The results indicate that increasing the number of edits does not significantly harm the model’s general capability (lines 3 - 5 and 7 - 9 in Table 9), demonstrating the potential scalability of our method. The continuous pre-training brings an inevitable modest side effect to the model (lines 2 - 3 and 6- 7 in Table 9).
@@ D.2 Side effect
Table 9: Side effect evaluation on MMLU [13].
D.3 Efficiency analysis
@@
Appendix E Societal impacts
The goal of our method is to conveniently and flexibly edit LLMs’ behavior for reasonable and benign purposes, such as providing new relevant information or fixing false or inaccurate responses. However, we caution readers that our method should not be used for any malicious or offensive purpose, including but not limited to, political authoritative facts, rumors, discriminatory statements, etc. It is worth reminding that the safe and responsible application of our method is really important. None of the offensive, toxic, or malicious editing should be allowed.

### v3

@@ B.1 Datasets
The dataset MQuAKE Zhong et al. (2023b) (Multi-hop Question Answering for Knowledge Editing) is constructed based on Wikidata and contains question answering instances that require 2-hop, 3-hop, and 4-hop reasoning. In the experiment, we use the latest version of the dataset 888”MQuAKE-CF-3k-v2.json” in https://github.com/princeton-nlp/MQuAKE, which fixes the knowledge conflict problem for the old version multi-edit subset, and report the accuracy for each query.
@@
DUNE Akyürek et al. (2023) is a benchmark designed for edits in natural language form. It evaluates the model’s capability of conducting natural language edits through four aspects: scientific reasoning, arithmetic reasoning, new information, and debiasing. As illustrated in Table 2 of Akyürek et al. (2023), the arithmetic reasoning edits do not follow natural language form as other subsets do and cannot represent a complete piece of instruction, therefore, we do not include it in our experiment.
WikiData-counterfact
The $\text{WikiData}_{counterfact}$ Cohen et al. (2024); Zhang et al. (2024a) collects triplets from top-viewed pages from Wikipedia and contains portability (ripple-effect Cohen et al. (2024)) instances to test whether the output to the input relevant to the edits is changed as well. Specifically, the portability evaluates the post-edited model from three aspects, including logical generalization, subject aliasing, and reasoning.
@@
The extended version of ZsRE Zhang et al. (2024a); Yao et al. (2023) is constructed based on the original ZsRE Levy et al. (2017), which is a dataset that focuses on the QA task. The extended version introduces a portability test Yao et al. (2023), including inverse relation, one-hop reasoning, and subject aliasing.
@@
COUNTERFACT Meng et al. (2023) is a dataset that concentrates on counterfactual information, which typically receives a lower prediction score than accurate facts. It constructs out-of-scope data by substituting the subject entity with a comparable description that has the same predicate.
B.2 Evaluation metrics
This section explains the evaluation metrics used in the extended ZsREYao et al. (2023); Zhang et al. (2024a) and $\text{Wiki}_{counterfact}$ Cohen et al. (2024); Zhang et al. (2024a). Generally, they adopt four metrics: reliability, generality, portability, and locality. Given an initial base model $f_{\theta}$, a post-edit model $f_{\theta^{\prime}}$, and a set of edit instances $(x_{t},y_{t})\in\{(x_{t},y_{t})\}$, the reliability is computed as the average accuracy of the edit cases:
@@ B.2 Evaluation metrics
For the extended ZsRE Yao et al. (2023); Zhang et al. (2024a) and $\text{Wiki}_{counterfact}$ Cohen et al. (2024); Zhang et al. (2024a), we follow the setting in the original paper and combine reliability and generality to the Edit Success rate.
B.3 Baseline implementation details
Unless otherwise specified, the baselines are implemented by using the EasyEdit framework Wang et al. (2023).
@@ B.3 Baseline implementation details
We follow the procedure implemented in previous work Meng et al. (2022, 2023); Yao et al. (2023); Zhang et al. (2024a) to fine-tune a specific layer from the model. We select layer 13 for Llama-3.2-1B and layer 27 for Qwen2.5-7B. For both models, we adopt the learning rate of $5e^{-4}$ and the number of optimization steps 25.
@@
For both models, we use LoRA Hu et al. (2022) to update the query and key projection matrix of the models, with rank set to 8, $\alpha$ set to 32, the dropout rate 0.1, and the learning rate $5e^{-3}$. The number of updating steps is set to 70 for Llama-3.2-1B and 60 for Qwen2.5-7B.
@@
ROME Meng et al. (2022) treats the FFN part of the LLMs as a key-value association and updates a pre-located layer by directly inserting an optimized key-value pair. We update the layer 5 for both Llama-3.2-1B and Qwen2.5-7B, and adopt 25 optimization steps for Llama-3.2-1B and 20 optimization steps for Qwen2.5-7B, with both learning rate $5e^{-5}$.
@@
R-ROME Gupta et al. (2024a) is another version of ROME Meng et al. (2022) with modified code implementation. We use the same hyperparameters as ROME.
@@
KN Dai et al. (2022) hypothesize that factual knowledge is stored in FFN memories and expressed by knowledge neurons. For both models, we use the threshold of 0.2 for knowledge attribution scores and 0.4 for the threshold of the prompts sharing percentage.
@@
GRACE Hartvigsen et al. (2022) adopts a discrete codebook to memorize the edits as key-value pairs. We set the location of the codebook layer 13 and 18 for Llama-3.2-1B and Qwen2.5-7B, respectively. Surprisingly, the $\epsilon$ value used in the original paper (1-3) seems insufficient for the complex editing experiments in this paper. Therefore, we increase it to 50. The number of optimization steps for the value vector is set to 100.
@@
IKE Zheng et al. (2023) maintains an explicit memory for edits and retrieves them via K-nearest neighbors. The retrieved edits are then used to construct demonstrations, which are then prefixed to the input to edit the behavior. In the experiments, we set $K=16$.
@@
MEND Mitchell et al. (2022a) trains an additional meta-network to predict a new rank-one update to the input gradient. In this paper, we train each model using ZsRE Levy et al. (2017) and COUNTERFACT Meng et al. (2023), and adopt the ZsRE-trained model to MQuAKE and the extended ZsRE and COUNTERFACT-trained model for $\text{Wiki}_{counterfact}$.
@@
SERAC Mitchell et al. (2022b) employs an explicit edit-instance memory, an additional trained scope classifier, and a trained counterfactual model. The scope classifier is responsible for determining whether an input is relevant to the edits in the memory. The input is fed to the counterfactual model once the input is deemed as relevant to memorized edits and the original model otherwise. We use the distilbert-base-cased Sanh et al. (2019) model as the scope classifier, and train it using the training set of ZsRE Levy et al. (2017) and COUNTERFACT Meng et al. (2023) from EasyEdit Wang et al. (2023). Following Akyürek et al. (2023), we use instruction-tuned models (Llama-3.2-1B-Instruct [^9] https://huggingface.co/meta-llama/Llama-3.2-1B-Instruct and Qwen2.5-0.5B [^10] https://huggingface.co/Qwen/Qwen2.5-0.5B-Instruct) for the counterfactual model.
@@
MEMIT Meng et al. (2023) is the extension of ROME Meng et al. (2022) that supports a batch of edits at a time. Unlike ROME, which only updates a single pre-located layer, MEMIT spreads the update to a set of identified layers. We apply changes to layers 4-8 for both models. Following the settings in the original paper, we set $\lambda$, the hyperparameter that balances the weighting of new and old associations, to $1.5\times 10^{4}$.
@@
EMMET Gupta et al. (2024b) is an unification of ROME Meng et al. (2022) and MEMIT Meng et al. (2023). Similar to ROME, we edit the layer 5 for both models, and set $\lambda=1e^{5}$.
@@
We adopt bge-base-en-v1.5 Xiao et al. (2023) as our retriever. For batch editing, we treat the corresponding batch number of edits as our corpus, and retrieve the 10 most relevant edits for each testing query.
@@
Appendix C Further Analysis
C.1 RAG vs. InComeS
@@ Appendix C Further Analysis
Method Model Single Editing Batch Editing 2-edits 3-edits 4-edits Avg 2-edits 3-edits 4-edits Avg Base Llama-3.2-1B 41.79 43.51 31.58 38.96 41.79 43.51 31.58 38.96 RAG 59.23 59.00 51.63 56.62 40.65 46.78 43.58 43.67 ICL 59.23 59.00 51.63 56.62 50.06 49.86 42.37 47.43 InComeS 71.19 72.17 72.62 71.99 53.93 52.79 52.73 53.15 Base Qwen2.5-7B 44.08 44.14 30.62 39.61 44.08 44.14 30.62 39.61 RAG 69.76 76.91 74.54 73.74 42.91 46.37 46.15 45.14 ICL 69.76 76.91 74.54 73.74 53.53 50.54 44.77 49.61 InComeS 66.46 71.24 76.54 71.41 55.13 53.48 47.91 52.17 Base Qwen3-8B-base 43.49 41.01 26.32 36.94 43.49 41.01 26.32 36.94 RAG 55.82 54.13 47.06 52.34 45.98 46.52 46.67 46.39 ICL 55.82 54.13 47.06 52.34 50.78 48.84 45.28 48.30 InComeS 56.36 58.71 59.93 58.33 53.42 50.90 50.36 51.56
Table 11: Results for RAG on MQuAKE Zhong et al. (2023b).
C.2 Side effect
We present a side effect analysis of our method in this section (Table 12). We test the editing side effect under three different numbers of edits (0, 0.1k, 1k) on MMLU Hendrycks et al. (2021) benchmark, which consists of 57 tasks across 4 domains, namely Social science, Humanities, STEM, and others. The results indicate that increasing the number of edits does not significantly harm the model’s general capability (lines 3 - 5 and 7 - 9 in Table 12), demonstrating the potential scalability of our method. The continuous pre-training brings an inevitable modest side effect to the model (lines 2 - 3 and 6- 7 in Table 12).
@@ C.2 Side effect
Table 12: Side effect evaluation on MMLU Hendrycks et al. (2021).
C.3 Efficiency analysis
@@
C.4 Applying loss on queries
C.5 Inclusion of zero gist
The motivation of including the zero-gist mechanism is to ensure that context-independent tokens can bypass the influence of the edit contexts. To assess the impact of zero-gist, we train a model without this mechanism and evaluate it on MQuAKE Zhong et al. (2023b) (see the “w/o zero-gist” line in Table 6). The results show a notable performance drop, suggesting that the cross-attention calculations may sometimes interfere with ordinary generation and our zero-gist strategy can mitigate this issue by allowing tokens to “attend to nothing”.
C.6 Information flow on tokens
C.7 Necessity of distillation
We verify the importance of the distillation loss in this section. The result (see line "w/o kl loss" and "w/o token weighting" in Table 6) shows a significant decrease when excluding any of the distillation components, indicating that the distillation plays an important role.
C.8 The edit success metric
As expected, Table 7 shows traditional editing methods such as fine-tuning and ROME exhibit high edit success rates; however, their portability scores generally lag behind the top-performing methods, highlighting a common limitation of these approaches. In contrast, ICL-based approaches that leverage the in-context learning capabilities of LLMs demonstrate superior performance in complex editing scenarios that require reasoning, owing to the enhanced context understanding of LLMs. We argue that the edit success metric may not be a comprehensive metric to evaluate the real capability of the post-edited model, as it can be cheated by the overfitting problem Zhang et al. (2025b), i.e., the model can assign disproportionately high probabilities to the edit target to get a high edit success rate. This may explain why the simple fine-tuning method shows an extremely high edit success rate but fails to maintain it on the porability. Therefore, we focus on complex settings, including multi-hop editing, natural language editing, and portability that aligns more with the real application scenarios. Our method demonstrates excellent performance in these settings, which aligns with our motivations described in Section 1.
C.9 Effectiveness of the method over model scale
We observe that the performance gain in Table 1 appears to diminish as the model scale increases. In this section, we conduct further experiments to find out whether this is true. The results in Table 13 show that the performance gain does not decrease when the model scale increases.
Method Model Single Editing Batch Editing 2-edits 3-edits 4-edits Avg 2-edits 3-edits 4-edits Avg Base Llama-3.2-1B 41.79 43.51 31.58 38.96 41.79 43.51 31.58 38.96 ICL 59.23 59.00 51.63 56.62 50.06 49.86 42.37 47.73 InComeS 71.19 72.17 72.62 71.99 53.93 52.79 52.73 53.15 Base Llama-3.1-8B 45.29 45.69 31.66 40.88 45.29 45.69 31.66 40.88 ICL 58.76 56.99 45.47 53.74 48.64 49.58 39.51 45.91 InComeS 74.39 78.25 75.24 75.96 56.28 56.55 51.90 54.91
Table 13: Results for different model scale Zhong et al. (2023b).
C.10 More results
To further verify the effectiveness of our method over recent model, we further test our method on Qwen3-8B-base. The results (Table 15 and Table 14) demonstrate that our method works well in recent models.
Method Model Single Editing Batch Editing 2-edits 3-edits 4-edits Avg 2-edits 3-edits 4-edits Avg Base Qwen3-8B-base 43.49 41.01 26.32 36.94 43.49 41.01 26.32 36.94 FT-M 51.32 50.39 45.22 48.98 48.11 44.24 41.39 44.58 LoRA 38.95 36.27 33.02 36.08 24.41 21.24 22.12 22.59 MEMIT 49.14 46.05 41.15 45.45 42.34 40.10 33.45 38.63 ICL 55.82 54.13 47.06 52.34 50.78 48.84 45.28 48.30 InComeS 56.36 58.71 59.93 58.33 53.42 50.90 50.36 51.56
Table 14: More results on mquake Zhong et al. (2023b).
Method Model Single Editing Batch Editing new info scientific de-biasing Avg new info scientific de-biasing Avg Base Qwen3-8B-base 82.99 86.62 31.07 66.89 82.99 86.62 31.07 66.89 FT-M 78.23 75.44 56.38 70.02 76.33 72.34 54.28 67.65 LoRA 71.23 69.99 43.23 61.48 73.13 64.59 42.22 59.98 MEMIT 81.29 80.45 49.65 70.46 80.99 78.53 45.66 68.39 ICL 83.06 81.12 45.45 69.88 83.35 82.76 30.34 65.48 InComeS 83.17 85.36 63.96 77.41 84.17 87.57 65.96 79.23
Table 15: More results on dune Akyürek et al. (2023).